Which of the points satisfy y-3 <_2x?
1. (0,0)
2. (1,5)
3. (-2,0)
A. 1, 2, and 3
B. 2 and 3 only
C. 1 and 2 only
D. 1 only

Answer by stanbon(48545)

Which of the points satisfy y-3 <_2x?
y <= 2x+3
----
Try (0,0): You get 0 <= 2*0+3 ;0 <= 3 ; that is true
Try (1,5): You get 5 <= 2*1+3 ;5 <= 5 ; that is true
Try (-2,0): You get 0 <= 2*-2+3 ; 0 <= -1 ; that is false
--------------------------------------------------
1. (0,0)
2. (1,5)
3. (-2,0)
------------------------------------------------------------
A. 1, 2, and 3
B. 2 and 3 only
C. 1 and 2 only
D. 1 only
========================
Ans: C
===============
Cheers,
Stan H.